OpenStudy (anonymous):

find the x and y intercepts of 6x-4y=20

OpenStudy (lgbasallote):

y int = -5 x int = 10/3 you know how that was got?

OpenStudy (anonymous):

no i dont

OpenStudy (lgbasallote):

y int -> let x = 0 6(0) - 4y = 20 0 - 4y = 20 -4y = 20 y = -5 x-int -> let y = 0 6x - 4(0) = 20 6x - 0 =20 6x = 20 x= 10/3 ^common factor of 2